OpenStudy (anonymous):

I have my answer, just want to know if the discriminant is right.. y=x^2-5x-8 y=x 2 real solutions and a discriminant of 68.

OpenStudy (anonymous):

Would the discriminant be 68 or 17?

OpenStudy (anonymous):

@Hero

hero (hero):

Would you mind showing your work ?

OpenStudy (anonymous):

Yeah one second. I used the quadratic formula -b+-sqrt b^2 - 4ac/2a

hero (hero):

Do you know what kind of problem this is you're working on?

OpenStudy (anonymous):

Yes. I just want to know if my answer is correct

OpenStudy (anonymous):

It says "use the discriminant to find the number of solutions for each system."

OpenStudy (freckles):

\[\text{ the disciminant of } y=ax^2+bx+c \text{ is } b^2-4ac\]

OpenStudy (anonymous):

I know

OpenStudy (freckles):

then how did you get two solutions?

OpenStudy (freckles):

discriminant will only be one number

OpenStudy (anonymous):

Because the answer it gives me is 6+-sqrt68/2 so I was wondering if the discriminant would be 68 or 17 once you simplify it.

OpenStudy (freckles):

so you didn't use the formula I wrote...

OpenStudy (anonymous):

I did!!

OpenStudy (freckles):

b^2-4ac is going to give you the discriminant

OpenStudy (freckles):

then how did you get + or - for the discriminant ?

OpenStudy (anonymous):

So it would be 68

OpenStudy (freckles):

D=b^2-4ac and you have a=1 b=-5 c=-8 just plug in D=(-5)^2-4(1)(-8) this shouldn't give you 68

OpenStudy (anonymous):

Yes it should. You set the two equations equal to each other and get x^2-6x-8=0

OpenStudy (freckles):

I see y=x^2-5x-8

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

Yes, but there is also the y=x

OpenStudy (freckles):

because I thought the question was find the discriminant of y=x^2-5x-8? Is that not the question?

OpenStudy (anonymous):

So once you put them equal to each other

OpenStudy (anonymous):

Read the entire thing. It is a system. You know with the little {

OpenStudy (anonymous):

{y=x^2-5x-8 {y=x

OpenStudy (freckles):

Oh I think you are trying to see how many solutions to the system there are I didn't understand the question because the question wasn't actually given in the title

OpenStudy (anonymous):

Yes....

OpenStudy (anonymous):

I just wanted to know if the 68 would give me 2 real solutions

OpenStudy (anonymous):

Is all I was asking

OpenStudy (freckles):

Yes replace y with x in the first equation \[x=x^2-5x-8 \\ 0=x^2-6x-8 \\ a=1\\ b=-6 \\ c=-8 \\ D=b^2-4ac=(-6)^2-4(1)(-8)=36+32=68>0\] yes 68>0 so that means you have two solutions

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Is what I was saying all along, but... :/

OpenStudy (freckles):

It might be good to post the full question in the title next time. It would be very helpful.

OpenStudy (anonymous):

Alright, alright.. Thank you nontheless

OpenStudy (freckles):

Like determine the number of solutions to the system :p

OpenStudy (anonymous):

*nonetheless

OpenStudy (freckles):

and np